The mass m and volume V of an object are

LaTeX: m=4.6\pm0.2

LaTeX: V=1.4\pm0.1

What is the density D = m/V of the object?

LaTeX: D_{av}=

LaTeX: \Delta D=
Answer #1

The formula for density is given by, D = m/V

Therefore, Dav = 4.6/1.4 = 3.3

The error can be found by using the formula, Dm

SD % 0.2.0.1 4.6 + 1.4

\Delta D = 0.115*3.3

\Delta D = 0.4

I hope i was able to answer your question. If you need anymore explanation, please ask in comments. I would be happy to help.

Please rate the solution if you are satisfied with the solution.